OpenStudy (anonymous):

What are the foci of the ellipse with the equation 16x^2+4y^2=64? Graph the ellipse. Please show all steps! I don't understand this. :/

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

c^2=a^2-b^2=16-4=12 c=√12 Foci: (0, 0±c)=(0,0±√12)=(0,-√12) and (0,√12) is this work correct, @Mertsj?

OpenStudy (mertsj):

You first need to divide both sides by 64 to get it into the right form. Did you do that?

OpenStudy (mertsj):

\[\frac{x^2}{4}+\frac{y^2}{16}=1\]

OpenStudy (anonymous):

Yeah I did that. Is writing it in the form I have above still correct?

OpenStudy (mertsj):

So x^2 is 16 and b^2 is 4 and c is sqrt12 or 2sqrt3 However, the big number is under the y so the major axis is vertical and the foci are c units above and c units below the center.

OpenStudy (mertsj):

So if you just simplify the sqrt 12, you are done. And you did a good job.

OpenStudy (whpalmer4):

Indeed!

OpenStudy (anonymous):

Thank you both! :)

OpenStudy (mertsj):

yw